# CSPCL: Category Semantic Prior Contrastive Learning for Deformable DETR-Based Prohibited Item Detectors

### v1

@@ 3.2 Category semantic prior contrastive learning (CSPCL)
where $\mathds{1}[k_{1}\neq k_{2}]\in\{0,1\}$ is an indicator function. It equals $1$ if $k_{1}\neq k_{2}$, and 00 in the other case. Additionally, $\mathcal{R}(k_{1},k_{2})$ is the repulsion factor, used to adjust the strength of the repulsive effect exerted by the category prototype on content queries from different categories. The higher the similarity between the prototype of the $k_{1}$-th category $\mathbf{p}_{i}^{k_{1}}$ and the prototype of the $k_{2}$-th category $\mathbf{p}_{j}^{k_{2}}$, the stronger the repulsion exerted by the prototype $\mathbf{p}_{i}^{k_{1}}$ on the content query $\mathbf{q}_{j}^{k_{2}}$. $\tau$ is the temperature coefficient, which controls the sensitivity of the IAR loss to inter-class prototypes similarity. As shown in Fig. 4, the larger the $\tau$, the more sensitive the IAR loss becomes to differences between prototypes. This means that the prototype only repels content queries belonging to similar categories while minimizing the influence on content queries from unrelated categories, which helps align the features between prototypes and queries. When $\tau=0$, $\mathcal{R}(k_{1},k_{2})$ becomes a constant, and the prototypes indiscriminately repel content queries from all categories.
@@ 3.3 Plug CSPCL into target decoder layers
The CSPCL mechanism can be inserted into the decoder of Deformable DETR-like models in a plug-and-play manner without increasing model complexity, as described in Algorithm 1. Given the number of classes $K$ in the dataset, the content queries $\mathbf{Q}_{c}$, localization results set $R$ and classification results set $C$ of all decoder layers set $L$, the class prototypes $\mathbf{P}$, and the set of target layers $\hat{L}$ where our CSPCL mechanism will be inserted. For each decoder layer $l$, the model’s classification and localization detection results are matched with the ground truth through the Hungarian matching mechanism [2] to establish a one-to-one correspondence. Then, we compute the built-in loss $\mathcal{L}_{BASE}$ of the model for the current layer $l$ based on the matched prediction result and ground truth pairs. Additionally, we determine whether the current layer belongs to the target layers set $\hat{L}$. If it does, we use the content queries from the current layer $\mathbf{Q}_{c}^{l}$, prototypes $\mathbf{P}$, and the number of categories $K$ to compute the CSP loss according to Eq. 8. This loss is then added to the model’s inherent loss to update the current layer’s loss value $\mathcal{L}^{l}$. Finally, we use the sum of the losses from each layer, denoted as $\mathcal{L}$, to update both the model parameters and the content queries of all layers.
Algorithm 1 Plug MMCL into the Decoder.
0:    Constant $K$; Set $L$, $\hat{L}$, $\mathbf{Q}_{c}$, $\mathbf{P}$, $\mathbf{R}$, $\mathbf{C}$, $\mathbf{G}$
@@
3.4 Analyze the effect of the CSPCL mechanism
Figure 5: The t-SNE visualization results of the content queries (points) and classifier weights (stars) from the first decoder layer. (a) and (b) show the visualization of content queries for DINO and DINO with the CSPCL mechanism, respectively. (c) and (d) show the visualization results of the classifier weights and content queries simultaneously reduced for the two models, DINO and DINO with the CSPCL mechanism, respectively.
Figure 6: The scatter plot and Kernel Density Estimation (KDE) joint distribution plot of prediction results of the final decoder layer. Blue and Orange represent the results of DINO and DINO with the CSPCL mechanism, respectively.
@@ 3.4 Analyze the effect of the CSPCL mechanism
To analyze the effectiveness of the CSPCL mechanism for content queries, we visualize their t-SNE [29] dimensionality reduction distribution results with classifier weights. As shown in Fig. 5(a) and Fig. 5(c), content queries of all categories of the original DINO are scattered and evenly distributed in the feature space, while also exhibiting significant differences from the category prototypes perceived by the classifier. This suggests that during the model training process, the content queries in the decoder fail to capture the semantic information responsible for classification, which is consistent with the conclusion drawn in Sec. 3.1. In contrast, Fig. 5(b) and Fig. 5(d) demonstrate that, under the CSPCL mechanism, intra-class content queries cluster together while maintaining necessary distances, whereas inter-class content queries repel each other. Additionally, the content queries of the same category show a clear correlation with the category prototypes perceived by the model, i.e., the classifier weights. This indicates that our method successfully aligns the category-specific content queries with the inherent feature distribution of contraband, thereby addressing the deficiency in Deformable DETR-like models.
@@
To analyze the effectiveness of the CSPCL mechanism for detection results, we visualize the IoU scores and classification scores of prediction results of DINO on the PIXray dataset, as shown in Fig. 6. We draw the scatter plot of prediction results whose classification scores and IoU scores are higher than 0.3, along with the Kernel Density Estimation (KDE) curves. Blue and orange represent the results of DINO and DINO with the CSPCL mechanism, respectively. The orange points are more concentrated and distributed further to the right compared to the blue points, indicating that under the CSPCL mechanism, the content queries capture more classification semantic prior information without losing localization features, improving the accuracy of classification results. This demonstrates that our CSPCL mechanism can help the model more effectively perceive and extract foreground information of specific classes from the overlapping features in X-ray images.
@@ 4.1 Experimental setup
Datasets and evaluation metrics. We conduct extensive experiments on two large-scale publicly available X-ray prohibited item datasets, OPIXray [33] and PIXray [23], which are widely used in the field of prohibited item detection. The PIXray dataset adopts the COCO [20] evaluation metric, AP, which measures the detector’s average precision across various IoU thresholds, providing a comprehensive assessment of overall detection performance. The OPIXray dataset adopts the VOC [8] evaluation metric, mAP, which is calculated as the mean average precision across all categories at an IoU threshold of 0.5.
@@
Implementation details. For fair, all training and testing are conducted on the same computer platform, equipped with an NVIDIA GeForce RTX 4090 GPU, an Intel Core i9-13900K CPU, 64GB of memory, running Windows 10, and using PyTorch 1.13.1. Furthermore, we utilize pre-trained ResNet-50 models from the official MMDetection [4] repository as the backbone, and all models are trained by the SGD optimizer with a learning rate of 0.01, a momentum of 0.9, and a weight decay of 0.1. In addition, all models are trained for 12 epochs, following the original training strategy, with an image size of $320\times 320$.
@@
Table 1: Ablation study about the CSPCL mechanism on DINO [37] on the PIXray [23] dataset. $L$ means all decoder layers. The superscript ’*’ represents the optimal value of the hyper-parameter.
@@
4.2 Ablation experiments
To obtain the optimal hyperparameters of CSP loss, we conduct extensive ablation experiments about integrating CSPCL into DINO on the PIXray dataset. As shown in Tab. 1, we first independently insert ITA loss at the 00-th decoder layer to adjust the truncation factor $\gamma$, with the optimal truncation factor $\gamma^{*}=5\times 10^{-3}$, achieving $65.4\%$ AP on the PIXray dataset. Similarly, as shown in Tab. 1, the independent insertion of IAR loss yields the optimal temperature factor $\tau^{*}=0.3$, achieving $65.6\%$ AP. Furthermore, Tab. 1 shows that when both ITA loss and IAR loss are used simultaneously, i.e., the full version of CSP loss, the model’s AP increases by $2.4\%$, which is higher than the individual uses of ITA loss and IAR loss. This demonstrates the compatibility of the two loss functions. Subsequently, we adjust the target layer set where the CSPCL is inserted. As shown in Tab. 1, the model achieves $67.2\%$ AP by aligning the content queries of all decoder layers with the category prototypes better than aligning any single layer. Notably, when aligning only one layer, shallower layers, such as layer 00, tend to perform better. This could be because deeper layers, such as layer $5$, involve content queries that more directly contribute to detection results, and indiscriminate alignment might disrupt the spatial features perceived by the model. Finally, as shown in Tab. 1 and Tab. 1, we adjust the parameters $\alpha$ and $\beta$ in Eq. 8 to control the contribution ratio of the two losses. The optimal values for $\alpha^{*}$ and $\beta^{*}$ are $1$ and $0.5$, respectively. Overall, the CSPCL mechanism with the optimal hyperparameters can help DINO achieve the highest AP score of $67.5\%$ on PIXray.
Method AP AP50 AP75 APS APM APL
DINO 64.3 86.5 71.0 19.3 48.9 73.9
DNIO + N-pair 66.0 (+1.7) 88.1 72.2 237 51.3 75.4
DINO + InfoNCE 65.4 (+1.1) 87.4 72.3 22.1 51.1 74.6
DINO + CSP (our) 67.5 (+3.2) 88.7 74.6 23.3 52.3 76.8
Table 2: Comparing classic contrastive losses with our CSP loss.
4.3 Generalization
Contrastive loss. To verify the generalization of the CSPCL mechanism for contrastive loss, we replace the proposed CSP loss with classic contrastive losses, such as N-pair loss and InfoNCE loss, to align content queries and prototypes of the same category, and evaluate them on DINO. As shown in Tab. 2, these losses still improve DINO’s AP score on the PIXray dataset by $1.7\%$ and $1.1\%$, respectively, but are less effective than the $3.2\%$ improvement achieved by CSP loss. From these results, we draw two conclusions: (1) The strategy of aligning content queries with category prototypes in the CSPCL mechanism is generally effective for prohibited item detection in X-ray images. (2) Thanks to gradient truncation and the adaptive repulsion mechanism, CSP loss is better suited for the multi-class multi-sample aligning task compared to the aforementioned classic contrastive losses.
Models and datasets. To verify the generalization of the CSPCL mechanism for models and datasets, we apply it to multiple representative Deformable DETR variants, including the basic Deformable DETR, the powerful DINO, RT-DETR, and the contraband-specialized AO-DETR on two prohibited item datasets PIXray and OPIXray. As shown in Tab. 3, the CSPCL mechanism improves the detection accuracy of Deformable DETR, RT-DETR, DINO, and AO-DETR on the PIXray dataset by $5.2\%$, $0.4\%$, $3.2\%$, and $0.8\%$ AP, respectively. Notably, CSPCL does not impose additional burdens on model complexity such as GFLOPs and PARAMs, nor does it reduce the model inference speed (FPS). Similarly, as shown in Tab. 4, CSPCL enhances the detection accuracy of the four models on the OPIXray dataset by $3.4\%$, $0.9\%$, $0.9\%$, and $0.8\%$ mAP, respectively. These experimental results demonstrate that the CSPCL mechanism can effectively improve the Deformable DETR-based models’ ability to handle overlapping detection tasks by aligning class prototypes and content queries.
Method CSPCL AP AP50 AP75 FPS PARAMs GFLOPs
Deformable DETR 36.6 66.4 37.1 60 52.14M 13.47
41.8 (+5.2) 72.3 44.5 60 52.14M 13.47
RT-DETR 61.4 84.0 68.5 64 42.81M 17.07
61.8 (+0.4) 84.3 68.7 64 42.81M 17.07
DINO 64.3 86.5 71.0 54 58.38M 26.89
67.5 (+3.2) 88.7 74.6 54 58.38M 26.89
AO-DETR 65.6 86.1 72.0 54 58.38M 26.89
66.4 (+0.8) 86.8 73.6 54 58.38M 26.89
Table 3: Generalization analysis of CSPCL mechanism on PIXray.
Method CSPCL mAP FO ST SC UT MU
Deformable DETR 65.6 68.1 30.1 88.9 66.0 75.0
69.0 (+3.4) 78.4 38.8 86.0 62.8 79.1
RT-DETR 69.2 75.6 30.5 88.4 66.7 84.9
70.1 (+0.9) 76.0 34.4 88.6 67.4 84.3
DINO 77.0 81.7 56.3 90.0 71.4 85.5
77.9 (+0.9) 82.8 56.0 89.9 74.2 86.7
AO-DETR 77.7 82.8 57.6 89.8 71.2 86.9
78.5 (+0.8) 84.3 57.7 90.2 73.2 87.1
Table 4: Generalization analysis of CSPCL strategy on OPIXray.

### v2

@@ 3.2 Category semantic prior contrastive learning (CSPCL)
where $\mathds{1}[k_{1}\neq k_{2}]\in\{0,1\}$ is an indicator function. It equals $1$ if $k_{1}\neq k_{2}$, and 0 in the other case. Additionally, $\mathcal{R}(k_{1},k_{2})$ is the repulsion factor, used to adjust the strength of the repulsive effect exerted by the category prototype on content queries from different categories. The higher the similarity between the prototype of the $k_{1}$-th category $\mathbf{p}_{i}^{k_{1}}$ and the prototype of the $k_{2}$-th category $\mathbf{p}_{j}^{k_{2}}$, the stronger the repulsion exerted by the prototype $\mathbf{p}_{i}^{k_{1}}$ on the content query $\mathbf{q}_{j}^{k_{2}}$. $\tau$ is the temperature coefficient, which controls the sensitivity of the IAR loss to inter-class prototype similarity. As shown in Figure˜3, the larger the $\tau$, the more sensitive the IAR loss becomes to differences between prototypes. This means that the prototype only repels content queries belonging to similar categories while minimizing the influence on content queries from unrelated categories, which helps align the features between prototypes and queries. When $\tau=0$, $\mathcal{R}(k_{1},k_{2})$ becomes a constant, and the prototypes indiscriminately repel content queries from all categories. Therefore, inter-class content queries can learn discriminative identifying features for similar categories, through our IAR loss with a satisfied $\tau$.
@@ 3.3 Plug CSPCL into target decoder layers
The CSPCL mechanism can be inserted into the decoder of Deformable DETR-like models in a plug-and-play manner without increasing model complexity, as described in Algorithm˜1. Given the number of classes $K$ in the dataset, the content queries $\mathbf{Q}_{c}$, localization results set $R$ and classification results set $C$ of all decoder layers set $L$, the class prototypes $\mathbf{P}$, and the set of target layers $\hat{L}$ where our CSPCL mechanism will be inserted. For each decoder layer $l$, the model’s classification and localization detection results are matched with the ground truth through the Hungarian matching mechanism [7] to establish a one-to-one correspondence. Then, we compute the built-in loss $\mathcal{L}_{BASE}$ of the model for the current layer $l$ based on the matched prediction result and ground truth pairs.
Additionally, we determine whether the current layer belongs to the target layers set $\hat{L}$. If it does, we use the content queries from the current layer $\mathbf{Q}_{c}^{l}$, prototypes $\mathbf{P}$, and the number of categories $K$ to compute the CSP loss according to Eq.˜5. This loss is then added to the model’s inherent loss to update the current layer’s loss value $\mathcal{L}^{l}$. Finally, we use the sum of the losses from each layer, denoted as $\mathcal{L}$, to update both the model parameters and the content queries of all layers.
Algorithm 1 Plug CSPCL into the Decoder.
0:Constant $K$; Set $L$, $\hat{L}$, $\mathbf{Q}_{c}$, $\mathbf{P}$, $\mathbf{R}$, $\mathbf{C}$, $\mathbf{G}$;
@@
3.4 Effect of the CSPCL mechanism
@@
To analyze the effectiveness of the CSPCL mechanism for content queries, we visualize their t-SNE [46] dimensionality reduction distribution results with classifier weights. As shown in Figure˜4(a) and Figure˜4(c), content queries of all categories of the original DINO are scattered and evenly distributed in the feature space, while also exhibiting significant differences from the category prototypes perceived by the classifier. This suggests that during the model training process, the content queries in the decoder fail to capture the semantic information responsible for classification, which is consistent with the conclusion drawn in Section˜3.1. In contrast, Figure˜4(b) and Figure˜4(d) demonstrate that, under the CSPCL mechanism, intra-class content queries cluster together while maintaining necessary distances, whereas inter-class content queries repel each other. Additionally, the content queries of the same category show a clear correlation with the category prototypes perceived by the model, i.e., the classifier weights. This indicates that our method successfully aligns the content queries with the inherent feature distribution of contraband, thereby addressing the deficiency in Deformable DETR-like models.
Figure 4: The t-SNE visualization results of the content queries (dots) and classifier weights (stars) from the first decoder layer. (a) and (b) show the visualization of content queries for DINO and DINO with the CSPCL mechanism, respectively. (c) and (d) show the visualization results of the classifier weights and content queries simultaneously reduced for the two models, DINO and DINO with the CSPCL mechanism, respectively.
@@
To analyze the effectiveness of the CSPCL mechanism for detection results, we visualize the IoU scores and classification scores of prediction results of DINO on the PIXray dataset, as shown in Figure˜5(a). We draw the scatter plot of prediction results whose classification scores and IoU scores are higher than 0.3, along with the Kernel Density Estimation (KDE) curves. Blue and orange represent the results of DINO and DINO with the CSPCL mechanism, respectively. The orange points are more concentrated and distributed further to the right compared to the blue points, indicating that under the CSPCL mechanism, the content queries capture more classification semantic prior information without losing localization features, improving the accuracy of classification results. This demonstrates that our CSPCL mechanism can help the model more effectively perceive and extract foreground information of specific classes from the overlapping features in X-ray images.
Figure 5: (a) The scatter plot and Kernel Density Estimation (KDE) joint distribution plot of prediction results of the final decoder layer. Blue and Orange represent the results of DINO and DINO with the CSPCL mechanism, respectively. (b) Visualization of deformable attention sampling points, reference points, and prediction results for corresponding group content queries in the last decoder layer. (i)–(iii) represent different groups. Each sampling point is shown as a filled circle, with color indicating its attention weight, and the reference point is marked by a green cross.
@@ 4.1 Experimental setup
Datasets and evaluation metrics. We conduct extensive experiments on four large-scale publicly available X-ray prohibited item datasets, OPIXray [2], PIXray [15], CLCXray [6], and PIDray [1], which are widely used in the field of prohibited item detection. The PIXray, CLCXray, and PIDray datasets adopt the COCO [47] evaluation metric, AP, which measures the detector’s average precision across various IoU thresholds, providing a comprehensive assessment of overall detection performance. The OPIXray dataset adopts the VOC [48] evaluation metric, mAP, which is calculated as the mean average precision across all categories at an IoU threshold of 0.5.
@@
Implementation details. For fair, all training and testing are conducted on the same computer platform, equipped with an NVIDIA GeForce RTX 4090 GPU, an Intel Core i9-13900K CPU, 64GB of memory, running Windows 10, and using PyTorch 1.13.1. Furthermore, we utilize pre-trained ResNet-50 models from the official MMDetection [49] repository as the backbone, and all models are trained by the SGD optimizer with a learning rate of 0.01, a momentum of 0.9, and a weight decay of 0.1. In addition, all models are trained for 12 epochs, following the original training strategy, with an image size of $320\times 320$ unless otherwise stated.
4.2 Generalization
Table 1: Comparing classic contrastive losses with our CSP loss under CSPCL mechanism.
Method AP AP50 AP75 APS APM APL
DINO 64.3 86.5 71.0 19.3 48.9 73.9
DNIO + N-pair 66.0 (+1.7) 88.1 72.2 23.7 51.3 75.4
DINO + InfoNCE 65.4 (+1.1) 87.4 72.3 22.1 51.1 74.6
DINO + CSP (our) 67.5 (+3.2) 88.7 74.6 23.3 52.3 76.8
Contrastive loss. To verify the generalization of the CSPCL mechanism for contrastive loss, we replace the proposed CSP loss with classic contrastive losses, such as N-pair loss and InfoNCE loss, to align content queries and prototypes of the same category, and evaluate them on DINO. As shown in Table˜1, these losses still improve DINO’s AP score on the PIXray dataset by $1.7\%$ and $1.1\%$, respectively, but are less effective than the $3.2\%$ improvement achieved by CSP loss. From these results, we draw two conclusions: (1) The strategy of aligning content queries with category prototypes in the CSPCL mechanism is generally effective for prohibited item detection in X-ray images; (2) Thanks to gradient truncation and the adaptive repulsion mechanism, CSP loss is better suited for the multi-class multi-sample aligning task compared to the aforementioned classic contrastive losses. More theoretical comparison analysis with other contrastive losses can be found in Appendix˜B.
Models and datasets. To verify the generalization of the CSPCL mechanism for models and datasets, we apply it to multiple representative Deformable DETR variants, including the basic Deformable DETR, the powerful DINO, RT-DETR, and the contraband-specialized AO-DETR on two prohibited item datasets PIXray and OPIXray. As shown in Table˜2, the CSPCL mechanism improves the detection accuracy of Deformable DETR, RT-DETR, DINO, and AO-DETR on the PIXray dataset by $5.2\%$, $0.4\%$, $3.2\%$, and $0.8\%$ AP, respectively. Notably, CSPCL does not impose additional burdens on model complexity such as GFLOPs and PARAMs, nor does it reduce the model inference speed (FPS). Similarly, CSPCL enhances the detection accuracy of the four models on the OPIXray dataset by $3.4\%$, $\%, 0.90.9\%$, and $0.8\%$ mAP, respectively. These experimental results demonstrate that the CSPCL mechanism can effectively improve the Deformable DETR-based models’ ability to handle overlapping detection tasks by aligning class prototypes and content queries.
Comparison with SOTA Models. To demonstrate the enhancement effect of CSPCL on the prohibited item detection model, we chose AO-DETR as the baseline and train it with CSPCL under the condition of $512\times 512$ image size, $14$ epochs, $240$ queries to challenge other SOTA models, including general detectors and prohibited item detectors, on the largest known dataset for the prohibited item detection dataset PIDray [1]. As shown in Table˜3, $\mathcal{C}$-AO-DETR (AO-DETR + CSPCL) obtains the highest AP over easy, hard, and hidden sets, while outperforms in terms of Image Size and FLOPs. The validation on the fine-grained dataset for cutters and liquid containers, CLCXray [6], is provided in Appendix˜A.
Table 2: Generalization analysis of CSPCL mechanism on PIXray and OPIXray.
Method CSPCL FPS PARAMs GFLOPs PIXray OPIXray
AP AP50 AP75 mAP FO ST SC UT MU
Deformable DETR 60 52.14M 13.47 36.6 66.4 37.1 65.6 68.1 30.1 88.9 66.0 75.0
60 52.14M 13.47 41.8 (+5.2) 72.3 44.5 69.0 (+3.4) 78.4 38.8 86.0 62.8 79.1
RT-DETR 64 42.81M 17.07 61.4 84.0 68.5 69.2 75.6 30.5 88.4 66.7 84.9
64 42.81M 17.07 61.8 (+0.4) 84.3 68.7 70.1 (+0.9) 76.0 34.4 88.6 67.4 84.3
DINO 54 58.38M 26.89 64.3 86.5 71.0 77.0 81.7 56.3 90.0 71.4 85.5
54 58.38M 26.89 67.5 (+3.2) 88.7 74.6 77.9 (+0.9) 82.8 56.0 89.9 74.2 86.7
AO-DETR 54 58.38M 26.89 65.6 86.1 72.0 77.7 82.8 57.6 89.8 71.2 86.9
54 58.38M 26.89 66.4 (+0.8) 86.8 73.6 78.5 (+0.8) 84.3 57.7 90.2 73.2 87.1
Table 3: Comparison with SOTA models on PIDray.
Method Backbone Image Size AP PARAMs FLOPs
easy hard hidden overall
Faster R-CNN [50] ResNeXt-101-32x4d $1333\times 800$ 70.0 64.5 49.7 61.4 60.2M 300.5G
Mask R-CNN[51] ResNeXt-101-32x4d $1333\times 800$ 77.7 70.0 53.7 67.0 101.9M 514.7G
FSAF [52] ResNeXt-101-64x4d $1333\times 800$ 71.0 61.9 50.7 61.2 94.0M 461.6G
FCOS [53] ResNeXt-101-64x4d $1333\times 800$ 72.9 63.4 51.1 62.5 89.6M 457.0G
ATSS [54] ResNet-101 $1333\times 800$ 71.0 64.5 55.4 63.6 51.1M 295.5G
VFNet [55] ResNet-101 - 70.3 62.8 48.6 60.6 32.74M -
TOOD [56] ResNet-101 - 68.5 63.8 48.9 60.4 32.04M -
UniFormer [57] UniFormer-B $1333\times 800$ 70.4 63.9 40.7 58.3 69.0M 399.0G
Deformable DETR [16] ResNet-50 $1333\times 800$ 72.1 65.7 46.2 61.3 41.0M 210.2G
DN-DETR [23] ResNet-50 $1333\times 800$ 76.8 71.0 55.6 67.8 47.0M 860.0G
SDANet [1] ResNet-101 $500\times 500$ 71.2 64.2 49.5 61.6 - -
ForkNet [39] ResNet-101 $500\times 500$ 75.0 66.9 58.6 66.8 - -
AO-DETR [10] Swin-L $512\times 512$ 81.8 72.6 55.7 70.0 229.0M 276.5G
$\mathcal{C}$-AO-DETR (ours) Swin-L $512\times 512$ 82.2 73.3 56.9 70.8 229.0M 276.5G
4.3 Ablation experiments
To obtain the optimal hyperparameters of CSP loss, we conduct extensive ablation experiments about integrating CSPCL into DINO on the PIXray dataset. As shown in Table˜4, we first independently insert ITA loss at the 0-th decoder layer to adjust the truncation factor $\gamma$, with the optimal truncation factor $\gamma^{*}=5\times 10^{-3}$, achieving $65.4\%$ AP on the PIXray dataset. Similarly, as shown in Table˜4, the independent insertion of IAR loss yields the optimal temperature factor $\tau^{*}=0.3$, achieving $65.6\%$ AP. Furthermore, Table˜4 shows that when both ITA loss and IAR loss are used simultaneously, i.e., the full version of CSP loss, the model’s AP increases by $2.4\%$, which is higher than the individual use of ITA loss and IAR loss. This demonstrates the compatibility of the two loss functions. Subsequently, we adjust the target layer set where the CSPCL is inserted. As shown in Table˜4, the model achieves $67.2\%$ AP by aligning the content queries of all decoder layers with the category prototypes better than aligning any single layer. Notably, when aligning only one layer, shallower layers, such as layer 0, tend to perform better. This could be because deeper layers, such as layer $5$, involve content queries that more directly contribute to detection results, and indiscriminate alignment might disrupt the spatial features perceived by the model. Finally, as shown in Table˜4 and Table˜4, we adjust the parameters $\alpha$ and $\beta$ in Eq.˜5 to control the contribution ratio of the two losses. The optimal values for $\alpha^{*}$ and $\beta^{*}$ are $1$ and $0.5$, respectively. Overall, the CSPCL mechanism with the optimal hyperparameters can help DINO achieve the highest AP score of $67.5\%$ on PIXray.
@@ 4.3 Ablation experiments
Table 4: Ablation study about the CSPCL mechanism on DINO [9] on the PIXray [15] dataset. $L$ means all decoder layers. The superscript ‘*’ represents the optimal value of the hyper-parameter.